Can anyone help out? I need to answer question 17 using the 5 step problem solving technique used in physics equationsCan anyone help out? I need to answer question 17
Answer #1

Step 1: ?t = 2.4 minutes (144 seconds), velocity from 0-1.2 minutes = 0.06m/s, velocity from 1.2-1.4 minutes = 12m/s

Step 2: Need to find the average speed over the 2.4 minute period.


Step 3: Find the total distance traveled and divide it by the total elapsed time because average speed = ?x/?t


Step 4: From 0-1.2, ?x1 = (0.06m/s)*(72s) = 4.32m. From 1.2-2.4, ?x2 = (12m/s)*(72s) = 864m. Total ?x = 868.32

average speed = ?x/?t. therefore average speed = 6.03m/s.

Step 5: Does this make sense? does the average speed times the total time = the total distance traveled? Yes it does.